OpenStudy (anonymous):

For the function defined by f(x)= x^2 x1.... identify the function that will be used when x=0?

OpenStudy (anonymous):

\[f(x)=\begin{cases}x^2&\text{for }x\le1\\2x+1&\text{for }x>1\end{cases}\] Which part of the function is defined for \(x=0\)?

OpenStudy (anonymous):

it says evaluate f(0)

OpenStudy (anonymous):

Right, I get that, but I'm asking you which piece of the piecewise function applies to that value of \(x\)?

OpenStudy (anonymous):

x^2 and 2x+1?

OpenStudy (anonymous):

Let me ask you a different way: Is 0 less than 1 or greater than 1?

OpenStudy (anonymous):

less than

OpenStudy (anonymous):

Okay, so which part of the given function is used for all numbers less than 1?

OpenStudy (anonymous):

x^2?

OpenStudy (anonymous):

Right, the rule for \(f(x)\) says for all values less than 1, it takes on \(f(x)=x^2\). So, since \(0<1\), you have \(f(0)=0^2=0\).

OpenStudy (anonymous):

ohh, i get it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw